how to get 50000.00 compounded 3% monthly in 20yrs
Answer by mananth(16946)

Principal P = x
Amount= 50000
years=n 20
compounded 12 times a year
Rate = 3.00 0.030
Amount = P*((n+r)/n)^n

50000 = x *( 1 + r/n )^ n


50000=P(1+0.030 /12)^12 *20

50000 =P( 1.00250 )^240

50000 /(1.00250)^240=P

P= 27461.14